# Hunting for a 17 MeV particle coupled to electrons

### v2

@@ VII Conclusions
In this work, we have tested the hypothesis of a new 17 MeV boson coupled to electrons, commonly referred to as $X_{17}$, by employing a set of observables complementary to those probed by PADME and other beam-dump experiments. We have shown that the $e^{+}e^{-}$ PADME excess is already in tension with the electron $g$-2, as well as exotic pion decay data from SINDRUM, although it remains marginally viable and testable.
We have analyzed a set of observables sensitive to the coupling of a light particle to electrons in the 1–100 MeV mass range, identifying regions of parameter space that are currently allowed and those soon to be probed. In particular, future measurements from Mu3e, PIONEER, and improved determinations of the electron $g$-2 will be able to definitively test the $X_{17}$ scenario in the near future.

### v3

@@ VII Conclusions
In this work, we have tested the hypothesis of a new 17 MeV boson coupled to electrons, commonly referred to as $X_{17}$, by employing a set of observables complementary to those probed by PADME and other beam-dump experiments. We have shown that the $e^{+}e^{-}$ PADME excess is already in tension with the electron gg-2, predicting a correlated signal with $|\Delta a_{e}|>10^{-12}$. The same holds for the SINDRUM experiment that effectively excludes the PADME excess, except in scenarios where $X_{17}$ is a spin-1 boson coupled to electrons and electron neutrinos with equal strength, or more generally to a conserved current, as in the case of a dark photon or $U(1)_{B-L}$.
We have analyzed a set of observables sensitive to the coupling of a light particle to electrons in the 1–100 MeV mass range, identifying regions of parameter space that are currently allowed and those soon to be probed. In particular, future measurements from Mu3e, PIONEER, and improved determinations of the electron gg-2 will be able to definitively test the $X_{17}$ scenario in the near future.
@@
Although we focused on the vector case motivated by the recent PADME analysis, we have also examined scalar, pseudoscalar, and axial-vector couplings, which may be relevant for other $X_{17}$ scenarios or in other contexts. As further data from PADME and related experiments becomes available, the framework explored here will remain a useful tool in scrutinizing the $X_{17}$ interpretation. Although we focused on the vector case motivated by the recent PADME analysis, we have also examined scalar, pseudoscalar, and axial-vector couplings, which may be relevant for other $X_{17}$ scenarios or in other contexts. As further data from PADME and related experiments becomes available, the framework explored here will remain a useful tool in scrutinizing the $X_{17}$ interpretation.